# Finely Stratified Rerandomization Designs

### v2

@@ 7.1 Asymptotically Exact Inference
$\widehat{v}_{1}$ =$n^{-1}\sum_{s\in\mathcal{S}^{\nu}_{n}}\frac{1}{a(s)-1}\sum_{i\not=j\in s}\widehat{m}_{i}\widehat{m}_{j}^{\prime}D_{i}D_{j}/p$
$\widehat{v}_{0}$ =$n^{-1}\sum_{s\in\mathcal{S}^{\nu}_{n}}\frac{1}{(k-a)(s)-1}\sum_{i\not=j\in s}\widehat{m}_{i}\widehat{m}_{j}^{\prime}(1-D_{i})(1-D_{j})/(1-p)$
Using these terms, construct the variance estimator
$\widehat{V}=\operatorname{Var}_{n}(\widehat{m}_{i})-v_{D}(\widehat{v}_{1}+\widehat{v}_{0}-\widehat{v}_{10}-\widehat{v}_{10}^{\prime}).$ (7.2)
We require a slight strengthening of our GMM assumptions 3.2.
@@
Under this condition, we can state our first inference result, showing consistent estimation of the asymptotic variance matrix in Theorem 6.3.
@@
By Theorem 6.3, $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{0})\Rightarrow N(0,V_{\phi}+V_{a}^{adj})$. Then the variance estimation result above allows for joint inference on $\theta_{0}$ using e.g. standard Wald-test or t-statistic based confidence regions.
7.2 Inference on the Finite Population Parameter
In this section, we provide asymptotically conservative inference on linear contrasts of the finite population parameter $c^{\prime}\theta_{n}$.
As noted above, the asymptotic variance $V_{a}^{adj}$ in Theorem 6.3 for estimating the finite population parameter $\theta_{n}$ is generically not identified. This happens because it depends on terms of the form $\operatorname{Var}(a|\psi)\propto\operatorname{Var}(g_{1}|\psi)+\operatorname{Var}(g_{0}|\psi)-2\operatorname{Cov}(g_{1},g_{0}|\psi)$, with $g_{d}=g(d,X,S(d),\theta_{0})$. However, $S(1)$ and $S(0)$ are never simultaneously observed (Neyman, (1990)), so $\operatorname{Cov}(g_{1},g_{0}|\psi)$ is generically not identified. We work with linear contrasts $c^{\prime}\theta_{n}$ since this allows us to tighten our upper bounds on the (non-identified) variance. To do so, let $\widehat{u}_{1}=E_{n}[\frac{D_{i}}{p}\widehat{m}_{i}\widehat{m}_{i}^{\prime}]-\widehat{v}_{1}$ and $\widehat{u}_{0}=E_{n}[\frac{1-D_{i}}{1-p}\widehat{m}_{i}\widehat{m}_{i}^{\prime}]-\widehat{v}_{0}$ using the estimator components above and consider the variance estimator
$\widehat{V}_{a}(c)=v_{D}([c^{\prime}\widehat{u}_{1}c]^{1/2}+[c^{\prime}\widehat{u}_{0}c]^{1/2})^{2}.$ (7.3)
By Theorem 6.3, we have $\sqrt{n}(c^{\prime}\widehat{\theta}_{adj}-c^{\prime}\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,c^{\prime}V_{a}^{adj}c)$. Our next result shows how to consistently estimate an upper bound on this asymptotic variance.
Theorem 7.3 (Inference).

### v3

@@ 7.2 Inference on the Finite Population Parameter
$\widehat{v}_{1}$ =$n^{-1}\sum_{s\in\mathcal{S}_{n}}\frac{1}{a(s)-1}\sum_{i\not=j\in s}\widehat{m}_{i}\widehat{m}_{j}^{\prime}D_{i}D_{j}/p$
$\widehat{v}_{0}$ =$n^{-1}\sum_{s\in\mathcal{S}_{n}}\frac{1}{(k-a)(s)-1}\sum_{i\not=j\in s}\widehat{m}_{i}\widehat{m}_{j}^{\prime}(1-D_{i})(1-D_{j})/(1-p)$
Collapsed Strata. If number of treated units $a(s)=1$ or $a(s)=k(s)-1$, as in matched pairs designs, the estimators above do not exist. In this case, we follow [^34] See Abadie and Imbens, (2008), Bai et al., (2021), Cytrynbaum, 2024b , Bai et al., 2024b for recent use of this method for inference on superpopulation parameters. In particular, Bai et al., (2021) showed asymptotic exactness of the collapsed strata method for matched pairs designs under the matching condition above. the method of collapsed strata (Hansen et al., (1953)), first agglomerating the original groups $s\in\mathcal{S}_{n}$ into larger groups satisfying $22\leq a(s)\leq k(s)-$ 2. For example, in a matched triples design with $p=1/3$, we agglomerate two triples into a larger group $s^{\prime}$ of $6$ units with $a(s^{\prime})=2$. To do so, for each $s\in\mathcal{S}_{n}$ define the centroid $\bar{\psi}_{s}=|s|^{-1}\sum_{i\in s}\psi_{i}$. Let $\nu:\mathcal{S}_{n}\to\mathcal{S}_{n}$ be a bijective matching between groups satisfying $\nu(s)\not=s$, $\nu^{2}=\operatorname{Id}$, and matching condition $\frac{1}{n}\sum_{s\in\mathcal{S}_{n}}|\bar{\psi}_{s}-\bar{\psi}_{\nu(s)}|_{2}^{2}=o_{p}(1)$. In practice, $\nu$ is obtained by matching the group centroids $\bar{\psi}_{s}$ into pairs using the Derigs, (1988) non-bipartite matching algorithm. Define $\mathcal{S}^{\nu}_{n}=\{s\cup\nu(s):s\in\mathcal{S}_{n}\}$ to be the enlarged groups. If $a(s)=1$ or $a(s)=k(s)-1$, we replace $\mathcal{S}_{n}$ with the larger groups $\mathcal{S}^{\nu}_{n}$ in the definitions of $\widehat{v}_{1}$ and $\widehat{v}_{0}$.
[…]
$\widehat{V}_{a}^{adj}(c)=v_{D}^{-1}([c^{\prime}\widehat{u}_{1}c]^{1/2}+[c^{\prime}\widehat{u}_{0}c]^{1/2})^{2}.$ (7.2)
To formalize this result we require a slight strengthening of GMM Assumption 3.2.
@@
Then the confidence interval $\widehat{C}_{fin}\equiv[c^{\prime}\widehat{\theta}_{adj}\pm z_{1-\alpha/2}\widehat{V}_{a}^{adj}(c)^{1/2}/\sqrt{n}]$ has coverage $P(c^{\prime}\theta_{n}\in\widehat{C}_{fin})\geq 1-\alpha-o(1)$ by Theorem 6.3 and Theorem 7.6.
The main result is stated for adjusted GMM estimation under stratified rerandomization, with ex-post adjustment to restore normality. For the case of pure stratification (no rerandomization) without adjustment, we can just set $w=0$ in the formulas above, obtaining a specialization $\widehat{V}_{a}(c)$ of $\widehat{V}_{a}^{adj}(c)$. We summarize this in a corollary:
Corollary 7.7 (Pure Stratification).